OpenStudy (anonymous):

Let \(x \in \mathbb{R}\). Show by applying the definition, that the absolute value function\( f(x):= |x|\) is not differentiable in \(x_{0} = 0\), although it is everywhere continuous. Is f differentiable for \(x_{0} \neq 0 \)?

OpenStudy (anonymous):

You have to distinguish two cases, x>0 and x <0. if x> 0, the derivative is 1 if x<0, the derivative is -1.

OpenStudy (anonymous):

Yeah F is differentiable for \[_{?}\neq0 with _{=0}\] because f(x)=\[\left| x \right|\] means f(x)=x when x is positive and -x when x is negative

OpenStudy (anonymous):

\[ \lim_{x\to 0} \frac { f(x) -f(0)}{x-0}=\lim_{x\to 0} \frac {|x|}{x}=\pm 1 \] depending if x >- or x <0 The limit does not exist and f is not differentiable at zero.

OpenStudy (anonymous):

But f is differentiable at any other x different from 0.

OpenStudy (anonymous):

ok Mr Elias it looks very good, this is our end solution right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you Mr @eliassaab very much and thank you too @pape

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you welcome